# Surrogate models for diffusion on graphs via sparse polynomials

### v2

@@ 2.4 Holomorphic regularity
Definition 2 (Holomorphic function).
@@
Definition 3 (Filled-in Berstein polyellipse).
@@
Assume that $f\in L_{\varrho}^{2}(\mathcal{U})$ can be extended in a holomorphic way to an open set $\mathcal{O}\subseteq\mathbb{C}^{d}$ such that $\mathcal{E}_{\rho}\subset\mathcal{O}$ for some $\boldsymbol{\rho}>\boldsymbol{1}$, i.e., there exists $\tilde{f}:\mathcal{O}\to\mathbb{C}$ holomorphic such that $\tilde{f}|_{\mathcal{U}}=f$. Then, the best $s$-term approximation error $\sigma_{s}(\boldsymbol{c})_{q}$ decays algebraically fast in $s$, with dimension-independent decay rate [1, Theorem 3.6]. Namely,
@@
for all $0<p\leq q\leq\infty$ and $s\in\mathbb{N}_{0}$, where $\boldsymbol{c}=(c_{\boldsymbol{\nu}})_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ is the infinite vector of coefficients of $f$, $C(d,p,\boldsymbol{\rho})>0$ is a constant depending on $d$, $p$ and $\boldsymbol{\rho}$ explicitly defined in [1, Theorem 3.6], and $L^{\infty}(\mathcal{\mathcal{E}}_{\boldsymbol{\rho}})$ is the space of essentially bounded functions $f:\mathcal{E}_{\boldsymbol{\rho}}\to\mathbb{C}$.
@@
with probability at least $1-\epsilon$; see [26, Theorem 6.1] and Appendix B. Similarly, denoting with $\hat{f}_{\mathsf{CS}}$ a solution to problem (10) respectively, we have
@@ 2.5 Community structure assumption for diffusion on graphs
The following Assumption, that will be used in the main result of this paper (Theorem 1), summarizes the requirements just discussed.
Assumption 2 (Holomorphic extension).
@@
We now provide a typical example of parametrization satisfying this assumption.
Example 1.
Given a parameter $\boldsymbol{y}=(y_{1},y_{2},...,y_{d})\in[-1,1]^{d}$, with $d$ as in (17), the matrix $\boldsymbol{C}(t,\boldsymbol{y})$ in equation (4) can be constructed using an affine parameterization, e.g.,
@@
where in the $(i,j)$-th block, $\mathds{1}\in\mathbb{R}^{\lvert\mathcal{C}_{i}\rvert\times\lvert\mathcal{C}_{j}\rvert}$ denotes a matrix of ones of suitable dimension, $\sigma:[K]\times[K]\mapsto[d]$ is a function that defines a unique ordering (e.g., lexicographic) on $\{(i,j)\in[K]^{2}:i\leq j\}$ and with $\sigma(i,j)=\sigma(j,i)$ for $i>j$. The functions $h_{1}(t),...,h_{d}(t)$ are continuous on $[0,T]$. Affine parameterizations are very popular in modeling parametric PDEs [1, 2] and represent a simple way to model diffusivity parameters ranging in some prescribed interval.
@@ 3 Theoretical guarantees
for a given node $v$, where $\boldsymbol{u}(\cdot,\boldsymbol{C}(\cdot,\boldsymbol{y}))$ is the unique solution to the problem (4) with $\boldsymbol{C}(t)=\boldsymbol{C}(t,\boldsymbol{y})$. As illustrated in Section 2.4, functions that can be extended in a holomorphic way have desirable properties with respect to the recoverability of their polynomial approximation; hence our goal is to show that there exists a holomorphic extension for $f$. We will use [23] and [1, Lemma 4.3] which states the following:
Lemma 1 (Holomorphic dependence; Volterra integral equations).
@@
Theorem 1 (Holomorphy of parametric graph diffusion).
@@
$B(\boldsymbol{u}_{0},T,\boldsymbol{M})=\|\boldsymbol{u}_{0}\|_{2}\cdot\exp\left(2\cdot\int_{0}^{T}\sup_{\boldsymbol{z}\in\mathcal{H}}\|\boldsymbol{M}(t,\boldsymbol{z})\|_{2\to 2}\,dt\right)<\infty$ (22)
and $\boldsymbol{M}$ is defined as in equation (5).
Proof.
@@
  • $\lVert\boldsymbol{h}(t,s,\boldsymbol{u},\boldsymbol{z})-\boldsymbol{h}(t,s,\boldsymbol{v},\boldsymbol{z})\rVert\leq L\lVert\boldsymbol{u}-\boldsymbol{v}\rVert$ for some Lipschitz constant $L>0$, $\forall t,s\in[0,T]$, $\forall\boldsymbol{u},\boldsymbol{v}\in\mathbb{C}^{|V|}$, $\forall\boldsymbol{z}\in\mathcal{K}$.
@@
Next, we want to show that $\boldsymbol{h}$ is continuous in $0,T]^{2}\times\mathbb{C}^{|V|}\times\mathcal{K}$, and holomorphic in $\mathring{\mathcal{K}}$. First recall that affine maps are holomorphic and that composition of holomorphic maps is holomorphic [27]. The map $\boldsymbol{z}\mapsto\tilde{\boldsymbol{C}}(t,\boldsymbol{z})$ is holomorphic by Assumption 2. For two generic matrices $\boldsymbol{A}$ and $\boldsymbol{B}$, the map $\boldsymbol{B}\mapsto\boldsymbol{B}\odot\boldsymbol{A}$ is linear in $\boldsymbol{B}$, therefore it is holomorphic. Furthermore, the composition of two holomorphic functions is holomorphic. Thus, the map $\boldsymbol{z}\mapsto\tilde{\boldsymbol{C}}(t,\boldsymbol{z})\odot\boldsymbol{W}(t)$ obtained from the composition of holomorphic maps is also holomorphic. Moreover, the map $\tilde{\boldsymbol{C}}\mapsto\delta_{ij}\sum_{k}\tilde{C}_{ik}W_{ik}$ is linear in $\tilde{\boldsymbol{C}}$, so it is holomorphic. Thus, the map $\boldsymbol{z}\mapsto\boldsymbol{D}(\tilde{\boldsymbol{C}}(\boldsymbol{z}))$, which is the composition of $\boldsymbol{z}\mapsto\tilde{\boldsymbol{C}}(\boldsymbol{z})$ and $\tilde{\boldsymbol{C}}\mapsto\delta_{ij}\sum_{k}\tilde{C}_{ik}W_{ik}$, is also holomorphic. Finally, the difference of two holomorphic functions is holomorphic. It then follows that $\boldsymbol{z}\mapsto\tilde{\boldsymbol{C}}(t,\boldsymbol{z})\odot\boldsymbol{W}(t)-\boldsymbol{D}(\tilde{\boldsymbol{C}}(\boldsymbol{z}))$ is holomorphic. This shows that $h$ is holomorphic.
@@
$\|\tilde{\boldsymbol{f}}\|_{L^{\infty}(\mathcal{H})}=\sup_{\boldsymbol{z}\in\mathcal{H}}\|\tilde{f}(\boldsymbol{z})\|_{2}=\sup_{\boldsymbol{z}\in\mathcal{H}}\|\tilde{\boldsymbol{u}}(T,\boldsymbol{z})\|_{2}.$ (23)
@@
Defining $\beta(t):=\sup_{\boldsymbol{z}\in\mathcal{H}}\|\tilde{\boldsymbol{M}}(t,\boldsymbol{z})\|_{2\to 2}$, we see that $\frac{\partial\boldsymbol{v}(t,\boldsymbol{z})}{\partial t}\leq\beta(t)\boldsymbol{v}(t,\boldsymbol{z})$, which, in turn, leads to $\tilde{\boldsymbol{v}}^{\prime}(t)\leq\beta(t)\tilde{\boldsymbol{v}}(t)$. We can now apply Gronwall’s lemma [29], obtaining
@@
which corresponds to (21). We conclude by showing that $\int_{0}^{T}\beta(t)dt<\infty$. Indeed, we have that $\beta$ is continuous in $t$, as $\|\cdot\|_{2\to 2}$ is continuous and $\tilde{\boldsymbol{M}}(t,\boldsymbol{z})$ is continuous. Therefore $\beta$ in bounded on $[0,T]$ and the integral is finite. This concludes the proof. ∎
@@ 3.1 Best $s$-term approximation rates
Based on Theorem 1, we can infer results about the convergence rate of the best $s$-term approximation error (recall Definition 1) for the solution map $f_{v}$ defined in equation (19). By combining [1, Theorem 3.6] and Theorem 1, we obtain the following corollary.
Corollary 1 (Best $s$-term approximation rates for parametric graph diffusion).
Let $T>0$ and let the parametric diffusivity matrix $\boldsymbol{C}:[0,T]\times\mathbb{R}^{|V|\times|V|}$ satisfy Assumption 2. Let the map $f_{v}:\mathcal{U}=[-1,1]^{d}\to\mathbb{C}$, be defined by $f_{v}(\boldsymbol{y})=u_{v}(T,\boldsymbol{y})$, where $v\in V$ and $\boldsymbol{u}$ is the solution of problem (4). Moreover, assume that the compact set $\mathcal{K}\subseteq\mathbb{C}^{d}$ of assumption 2 is such that $\mathcal{E}_{\boldsymbol{\rho}}\subset\mathring{\mathcal{K}}$ for some $\boldsymbol{\rho}>\boldsymbol{1}$; let $B(u_{0},T,\boldsymbol{M)}$ be the bound in equation (22) and let $\boldsymbol{c}=(c_{\boldsymbol{\nu}})_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ be the sequence of coefficient of $f_{v}$ with respect to either the Chebyshev or Legendre basis, i.e., $f_{v}=\sum_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}c_{\boldsymbol{\nu}}\psi_{\boldsymbol{\nu}}$. Then, $\forall s\in\mathbb{N}_{0}$, and $0<p\leq q<\infty$ it holds
$\sigma_{s}(\boldsymbol{c})_{q}\leq\frac{B(u_{0},T,\boldsymbol{M})\cdot C(p,\boldsymbol{\rho},d)}{(s+1)^{\frac{1}{p}-\frac{1}{q}}}.$
This result implies that the best $s$-term approximation error $\sigma_{s}(\boldsymbol{c})_{q}$ with respect to Legendre or Chebyshev polynomials decays algebraically fast in $s$. In particular, $f_{v}$ defined in equation (19) admits a holomorphic extension to some compact $\mathcal{K}$, therefore the best $s$-term approximation error is bounded from above as in equation (15). The term $\lVert f_{v}\rVert_{L^{\infty}(\mathcal{E}_{\boldsymbol{\rho}})}$ in equation (15) is bounded as in Theorem 1, with $\mathcal{H}=\mathcal{E}_{\boldsymbol{\rho}}$.

### v3

@@ 2.4 Holomorphic regularity
Definition 2 (Holomorphic function)
@@
Definition 3 (Filled-in Berstein polyellipse)
@@
Assume that $f\in L_{\varrho}^{2}(\mathcal{U})$ can be extended in a holomorphic way to an open set $\mathcal{O}\subseteq\mathbb{C}^{d}$ such that $\mathcal{E}_{\rho}\subset\mathcal{O}$ for some $\boldsymbol{\rho}>\boldsymbol{1}$, i.e., there exists $\tilde{f}:\mathcal{O}\to\mathbb{C}$ holomorphic such that $\tilde{f}|_{\mathcal{U}}=f$. Then, the best $s$-term approximation error $\sigma_{s}(\boldsymbol{c})_{q}$ decays algebraically fast in $s$, with dimension-independent decay rate (sparsepoly, , Theorem 3.6). Namely,
@@
for all $0<p\leq q\leq\infty$ and $s\in\mathbb{N}_{0}$, where $\boldsymbol{c}=(c_{\boldsymbol{\nu}})_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ is the infinite vector of coefficients of $f$, $C(d,p,\boldsymbol{\rho})>0$ is a constant depending on $d$, $p$ and $\boldsymbol{\rho}$ explicitly defined in (sparsepoly, , Theorem 3.6), and $L^{\infty}(\mathcal{\mathcal{E}}_{\boldsymbol{\rho}})$ is the space of essentially bounded functions $f:\mathcal{E}_{\boldsymbol{\rho}}\to\mathbb{C}$.
@@
with probability at least $1-\epsilon$; see (adcock2022monte, , Theorem 6.1) and Appendix B. Similarly, denoting with $\hat{f}_{\mathsf{CS}}$ a solution to problem (10) respectively, we have
@@ 2.5 Community structure assumption for diffusion on graphs
Remark 1(Interpretation of the parameter $\boldsymbol{y}$)
Throughout the paper we will interpret $\boldsymbol{y}\in[-1,1]^{d}$ as a deterministic vector defining the parametric diffusivity matrix $\boldsymbol{y}\mapsto\boldsymbol{C}(\cdot,\boldsymbol{y})$. However, it is worth pointing out that in applications involving uncertainty quantification smith2024uncertainty the parameter $\boldsymbol{y}$ is typically a random vector modelling volatility in the diffusivity matrix $\boldsymbol{C}(\cdot,\boldsymbol{y})$ (which, in that scenario, would also be a random variable).
@@
The following Assumption, that will be used in the main result of this paper (Theorem 3.1), summarizes the requirements just discussed.
Assumption 2 (Holomorphic extension)
@@
Whereas the presence of a community structure in the graph and the continuity of the diffusivity matrix with respect to the time variable are intrinsic features of the problem, the specific form of the parametrization $\boldsymbol{y}\mapsto\boldsymbol{C}(\cdot,\boldsymbol{y})$ can be designed by the user. Hence, Assumption 2 should be thought of as a guiding principle to be followed while parametrizing the family of problems of interest. We now provide a typical example of parametrization satisfying this assumption. In the numerical experiments of Section 4 only parametric diffusivity matrices of this form will be considered.
Example 1
Given a parameter $\boldsymbol{y}=(y_{1},y_{2},...,y_{d})\in[-1,1]^{d}$, with $d$ as in (17), the matrix $\boldsymbol{C}(t,\boldsymbol{y})$ in equation (4) can be constructed using an affine parametrization, e.g.,
@@
where in the $(i,j)$-th block, $\mathds{1}\in\mathbb{R}^{\lvert\mathcal{C}_{i}\rvert\times\lvert\mathcal{C}_{j}\rvert}$ denotes a matrix of ones of suitable dimension, $\sigma:[K]\times[K]\mapsto[d]$ is a function that defines a unique ordering (e.g., lexicographic) on $\{(i,j)\in[K]^{2}:i\leq j\}$ and with $\sigma(i,j)=\sigma(j,i)$ for $i>j$. The functions $h_{1}(t),...,h_{d}(t)$ are continuous on $[0,T]$. Affine parametrizations are very popular in modeling parametric PDEs sparsepoly ; cohen2015approximation and represent a simple way to model diffusivity parameters ranging in some prescribed interval.
@@ 3 Theoretical guarantees
for a given node $v$, where $\boldsymbol{u}(\cdot,\boldsymbol{C}(\cdot,\boldsymbol{y}))$ is the unique solution to the problem (4) with $\boldsymbol{C}(t)=\boldsymbol{C}(t,\boldsymbol{y})$. As illustrated in Section 2.4, functions that can be extended in a holomorphic way have desirable properties with respect to the recoverability of their polynomial approximation; hence our goal is to show that there exists a holomorphic extension for $f$. We will use walter and (sparsepoly, , Lemma 4.3) which states the following:
Lemma 1 (Holomorphic dependence; Volterra integral equations)
@@
Remark 2(Choice of the vector norm $\|\cdot\|$)
The Lipschitz condition (20) of Lemma 1 depends on the choice of vector norm $\|\cdot\|$. We observe that this choice has no practical implications on the conclusion of the lemma. In the proof of Theorem 3.1, which will crucially rely on Lemma 1, we will choose $\|\cdot\|=\|\cdot\|_{\infty}$ because this simplifies computations, but other equally valid choices could be made.
@@
Theorem 3.1(Holomorphy of parametric graph diffusion)
@@
$B(\boldsymbol{u}_{0},T,\tilde{\boldsymbol{M}},\mathcal{H})=\|\boldsymbol{u}_{0}\|_{2}\cdot\exp\left(2\cdot\int_{0}^{T}\sup_{\boldsymbol{z}\in\mathcal{H}}\|\tilde{\boldsymbol{M}}(t,\boldsymbol{z})\|_{2\to 2}\,dt\right)<\infty,$ (22)
where $\tilde{\boldsymbol{M}}(t,\cdot):=\tilde{\boldsymbol{C}}(t,\cdot)\odot\boldsymbol{W}(t)-\boldsymbol{D}(\tilde{\boldsymbol{C}}(t,\cdot),\boldsymbol{W}(t))$ is a holomorphic extension of the function $\boldsymbol{M}(t,\cdot)$ defined in (5) to $\mathcal{K}$ for every $t\in[0,T]$, and $\tilde{\boldsymbol{C}}$ is as in Assumption 2.
Proof
@@
  • $\lVert\boldsymbol{h}(t,s,\boldsymbol{u},\boldsymbol{z})-\boldsymbol{h}(t,s,\boldsymbol{v},\boldsymbol{z})\rVert\leq L\lVert\boldsymbol{u}-\boldsymbol{v}\rVert$ for some Lipschitz constant $L>0$, $\forall t,s\in[0,T]$, $\forall\boldsymbol{u},\boldsymbol{v}\in\mathbb{C}^{|V|}$, $\forall\boldsymbol{z}\in\mathcal{K}$. Here we will verify this condition for $\|\cdot\|=\|\cdot\|_{\infty}$.
@@
Next, we want to show that $\boldsymbol{h}$ is continuous in $0,T]^{2}\times\mathbb{C}^{|V|}\times\mathcal{K}$, and holomorphic in $\mathring{\mathcal{K}}$. First recall that affine maps are holomorphic and that composition of holomorphic maps is holomorphic henrici1993applied . The map $\boldsymbol{z}\mapsto\tilde{\boldsymbol{C}}(t,\boldsymbol{z})$ is holomorphic by Assumption 2. For two generic matrices $\boldsymbol{A}$ and $\boldsymbol{B}$, the map $\boldsymbol{B}\mapsto\boldsymbol{B}\odot\boldsymbol{A}$ is linear in $\boldsymbol{B}$, therefore it is holomorphic. Furthermore, the composition of two holomorphic functions is holomorphic. Thus, the map $\boldsymbol{z}\mapsto\tilde{\boldsymbol{C}}(t,\boldsymbol{z})\odot\boldsymbol{W}(t)$ obtained from the composition of holomorphic maps is also holomorphic. Moreover, the map $\tilde{\boldsymbol{C}}\mapsto\delta_{ij}\sum_{k}\tilde{C}_{ik}W_{ik}$ is linear in $\tilde{\boldsymbol{C}}$, so it is holomorphic. Thus, the map $\boldsymbol{z}\mapsto\boldsymbol{D}(\tilde{\boldsymbol{C}}(\boldsymbol{z}))$, which is the composition of $\boldsymbol{z}\mapsto\tilde{\boldsymbol{C}}(\boldsymbol{z})$ and $\tilde{\boldsymbol{C}}\mapsto\delta_{ij}\sum_{k}\tilde{C}_{ik}W_{ik}$, is also holomorphic. Finally, the difference of two holomorphic functions is holomorphic. It then follows that $\boldsymbol{z}\mapsto\tilde{\boldsymbol{C}}(t,\boldsymbol{z})\odot\boldsymbol{W}(t)-\boldsymbol{D}(\tilde{\boldsymbol{C}}(\boldsymbol{z}))$ is holomorphic. This shows that $h$ is holomorphic.
@@
$\|\tilde{\boldsymbol{f}}\|_{L^{\infty}(\mathcal{H})}=\sup_{\boldsymbol{z}\in\mathcal{H}}\|\tilde{\boldsymbol{f}}(\boldsymbol{z})\|_{2}=\sup_{\boldsymbol{z}\in\mathcal{H}}\|\tilde{\boldsymbol{u}}(T,\boldsymbol{z})\|_{2}.$ (23)
@@
Defining $\beta(t):=\sup_{\boldsymbol{z}\in\mathcal{H}}\|\tilde{\boldsymbol{M}}(t,\boldsymbol{z})\|_{2\to 2}$, we see that $\frac{\partial\boldsymbol{v}(t,\boldsymbol{z})}{\partial t}\leq\beta(t)\boldsymbol{v}(t,\boldsymbol{z})$, which, in turn, leads to $\tilde{\boldsymbol{v}}^{\prime}(t)\leq\beta(t)\tilde{\boldsymbol{v}}(t)$. We can now apply Gronwall’s lemma gronwall1919note , obtaining
@@
which corresponds to (21). We conclude by showing that $\int_{0}^{T}\beta(t)dt<\infty$. Indeed, we have that $\beta$ is continuous in $t$, as $\|\cdot\|_{2\to 2}$ is continuous and $\tilde{\boldsymbol{M}}(t,\boldsymbol{z})$ is continuous. Therefore $\beta$ in bounded on $[0,T]$ and the integral is finite. This concludes the proof.
@@ 3.1 Best $s$-term approximation rates
Based on Theorem 3.1, we can infer results about the convergence rate of the best $s$-term approximation error (recall Definition 1) for the solution map $f_{v}$ defined in equation (19). In fact, by using the holomorphy of $f_{v}$ one can show that its coefficients with respect to Legendre or Chebyshev orthogonal polynomials exhibit an exponential decay. In turn, an application of Stechkin’s inequality (see, e.g., (dung2018hyperbolic, , Section 7.4)) allows one to turn an exponential coefficient decay bound into a super-algebraic best $s$-term approximation rate estimate. Therefore, this shows that holomorphic regularity implies approximate sparsity (or compressibility) with respect to these orthogonal polynomial families. An in-depth survey of this theory can be found in (sparsepoly, , Chapter 3). Here, by combining (sparsepoly, , Theorem 3.6) (see also cohen2015approximation ; cohen2011analytic ) and Theorem 3.1, we obtain the following corollary.
Corollary 1 (Best $s$-term approximation rates for parametric graph diffusion)
Let $T>0$ and let the parametric diffusivity matrix $\boldsymbol{C}:[0,T]\times\mathbb{R}^{|V|\times|V|}$ satisfy Assumption 2. Let the map $f_{v}:\mathcal{U}=[-1,1]^{d}\to\mathbb{C}$, be defined by $f_{v}(\boldsymbol{y})=u_{v}(T,\boldsymbol{y})$, where $v\in V$ and $\boldsymbol{u}$ is the solution of problem (4). Moreover, assume that the compact set $\mathcal{K}\subseteq\mathbb{C}^{d}$ of assumption 2 is such that $\mathcal{E}_{\boldsymbol{\rho}}\subset\mathring{\mathcal{K}}$ for some $\boldsymbol{\rho}>\boldsymbol{1}$; let $B(\boldsymbol{u}_{0},T,\tilde{\boldsymbol{M}},\mathcal{E}_{\boldsymbol{\rho}})$ be the bound in equation (22) and let $\boldsymbol{c}=(c_{\boldsymbol{\nu}})_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ be the sequence of coefficient of $f_{v}$ with respect to either the Chebyshev or Legendre basis, i.e., $f_{v}=\sum_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}c_{\boldsymbol{\nu}}\psi_{\boldsymbol{\nu}}$. Then, $\forall s\in\mathbb{N}_{0}$, and $0<p\leq q<\infty$ it holds
$\sigma_{s}(\boldsymbol{c})_{q}\leq\frac{B(\boldsymbol{u}_{0},T,\tilde{\boldsymbol{M}},\mathcal{E}_{\boldsymbol{\rho}})\cdot C(p,\boldsymbol{\rho},d)}{(s+1)^{\frac{1}{p}-\frac{1}{q}}},$
where the constant $C(p,\boldsymbol{\rho},d)>0$ is as in (sparsepoly, , Theorem 3.6).
This result implies that the best $s$-term approximation error $\sigma_{s}(\boldsymbol{c})_{q}$ with respect to Legendre or Chebyshev polynomials decays algebraically fast in $s$. In particular, $f_{v}$ defined in equation (19) admits a holomorphic extension to some compact $\mathcal{K}$, therefore the best $s$-term approximation error is bounded from above as in equation (15). The term $\lVert f_{v}\rVert_{L^{\infty}(\mathcal{E}_{\boldsymbol{\rho}})}$ in equation (15) is bounded as in Theorem 3.1, with $\mathcal{H}=\mathcal{E}_{\boldsymbol{\rho}}$. A numerical illustration of Corollary 1 will be provided in Figures 2 and 11.
The main takeaway of Corollary 1 is that the map $f_{v}$ is compressible with respect to Legendre and Chebyshev polynomials. This, in turn, implies that sparse polynomial methods are a suitable option to approximate it. The next section will provide convergence guarantees for such methods.